# Modeling Language as a Sequence of Thoughts

### v1

@@ 3.3 Training schedules
Sentence-stream curriculum (chunk-size curriculum).
Retaining computation graphs through memory writes causes the effective backpropagation depth to grow with the number of sentences processed continuously. Early in training, sentence representations are largely uninformative; long sentence streams therefore increase compute and optimization difficulty without providing useful long-range credit assignment. We address this with a curriculum over the maximum sentence-stream length used to slice training documents: we begin with shorter streams and periodically increase and re-chunk the training split (validation and test sets are never split into sentence streams). In our experiments with memory capacity , we start at sentences and increase by every 5 epochs. This expands the effective dependency span of optimization from 750 (average sentence length of ) to 2000 tokens over the course of training, while keeping early optimization focused on within-sentence token modeling and short-range memory use.
@@
In TG, end-of-sentence markers (<EOS>) are more frequent than lexical tokens, as they appear at the end of every sentence and are often comparatively easy to predict (strongly signaled by punctuation and syntax). To mitigate this frequency and easy-label imbalance, we apply a frequency-aware reweighting of the token-level loss that down-weights <EOS> targets after an initial warm-up epoch (weight 1.0 for epoch 1, then 0.05 thereafter) [61, 62]. This affects the training loss the model receives; separately, special-token targets (<EOS>/<EOD>) are excluded from reported perplexities and checkpoint selection, as described in §4. We also use a warm-in schedule for stochastic regularization detailed in Appendix C.2.
@@ 4 Results
We evaluate Thought Gestalt (TG) through a series of experiments designed to measure data efficiency and representational robustness. First, we analyze TG’s scaling behavior by measuring how loss improves with training data and model size relative to a standard GPT-2 baseline, following the empirical scaling-law framework of Kaplan et al. [63]. Second, we benchmark TG’s data scaling against three architectural baselines that isolate key design concepts: (1) GPT-2 with Sentence Boundary Bias, which induces structural bias by retaining sentence start/end tokens (<BOS>, <EOS>) in the token stream; (2) Fixed Token-Span Recurrence, which replaces TG’s sentence-level processing with fixed-length token windows similar to prior recurrent work [49, 48]; and (3) GPT-2 + Gist Masking, where tokens pay causal attention to prior tokens only within a sentence and to preceding sentences through compressed "gist" tokens [38]. This attention mechanism is similar to TG but utilizes only in-context attention masks without recurrence or external memory. Third, we assess relational direction generalization, which evaluates robustness to the in-context reversal curse given a controlled father–son probe [20, 21]. Finally, we conduct design ablations to identify how specific components, such as gradient flow through memory, drive performance gains.
Experimental setup.
@@ 4.1 Scaling Efficiency
We quantify the data and parameter efficiency of Thought Gestalt (TG) using the empirical scaling-law framework of Kaplan et al. [63]. Following this work, we treat as the held-out cross-entropy loss in nats/token and as the number of non-embedding model parameters. When other factors are not the limiting bottleneck, Kaplan et al. observed approximate power-law relationships and , which appear as straight lines in log–log space. In this regime, architectural improvements typically manifest as a downward shift of the curve (lower loss at fixed resources). We first fit these laws to TG and a matched GPT-2 baseline trained under identical optimization and evaluation protocols described in sec. 4 to compare their data and parameter efficiency.
@@
Figure 2(a) shows that TG achieves lower test perplexity at every dataset size. Fitting a power law to the test loss yields similar scaling exponents for the two models (TG: ; GPT-2: ), suggesting that over this range the gain is primarily due to an intercept shift. Across our tested range this corresponds to a – reduction in perplexity (e.g., at M tokens, TG reaches 23.2 test perplexity versus 24.0 for GPT-2; see Appendix B.1 for other perplexities). To express the improvement as data efficiency, we compute an effective data multiplier , where is TG’s achieved test loss at dataset size , and is the number of training tokens GPT-2 would need to reach loss according to the fitted GPT-2 scaling curve. We find –, meaning GPT-2 requires about 5–8% more training tokens to match TG’s loss; supporting the hypotheses that learning latent representations at more abstract levels than tokens can lead to more sample efficient learning.
@@ 4.2 Other Baselines
TG leverages sentence boundaries as a structural bias to process information in semantically coherent units. Here, we evaluate whether a standard Transformer benefits simply from making this structure explicit in the token stream. To do this, we take the sentence-segmented output of the TG data pipeline (§3.1) and flatten it into a continuous sequence for a standard GPT-2 model. Crucially, we retain the <BOS> and <EOS> markers to explicitly delimit boundaries, while removing the padding tokens required for TG’s fixed-size sentence tensors. In this setup, <BOS> predicts the sentence’s first lexical token, the last lexical token predicts <EOS>, and <EOS> predicts the <BOS> of the subsequent sentence. The model architecture and attention mechanism remain identical to the GPT-2 baseline. To avoid making the comparison hinge on predicting frequent boundary events, we compute validation loss for checkpoint selection and report perplexity over lexical tokens only, excluding special-token label positions, matching the TG evaluation protocol (§4). Figure 5(left) shows that inducing sentence boundary bias improves GPT-2 performance across all dataset sizes, with the most significant gains in the low-data regime (test PPL decreases at 20M; see Table 3). However, the improvement shrinks with scale (e.g., at 50M), and TG surpasses this baseline once more data is available (TG: vs. at 50M). This indicates that boundary tokens alone provide a helpful structural inductive bias for token LMs, but this is not sufficient to replicate the efficiency gains derived from TG’s learning and reuse of contextualized sentence-level latent states.
@@
TG’s attention pattern—full causal token interaction within a sentence, and access to earlier content only through compact summary vectors—is similar to gisting: tokens attend to their current segment, and later tokens can only access prior segments through a small set of “gist” representations learned via attention masking in a standard Transformer [38]. To test whether this attention-distribution mechanism can reproduce TG’s gains without recurrence or an external memory, we implement a GPT-2 baseline that (i) inserts <BOS>/<EOS> boundaries in the token stream and (ii) applies an additive attention-bias mask that restricts each token to attend causally within its own sentence, while accessing previous sentences only via each sentence’s last token, <EOS>, acting as a “gist” token. This matches TG’s high-level attention connectivity, but removes recurrence, which has a key stabilizing effect: TG exposes fully contextualized sentence vectors (extracted after within-sentence processing and memory reads) to every layer via cross-attention, whereas GPT-2 + gist masking forces layers to read from in-context compression tokens whose representations are being formed in parallel with all other tokens (The effect of placing compression vectors in-context (as in gisting) versus in an external memory is detailed in §4.4). Empirically, the gist-masking baseline performs worse across all data scales (test PPL 31.9 at 50M; Table 5), far behind both GPT-2 (24.0 at 50M) and TG (23.2 at 50M), as seen in Figure 5 (right). These results show that when compressed states are not reliably contextualized, restricting direct access to prior tokens can amplify errors rather than improve abstraction.
@@
We evaluate the robustness of TG and GPT-2 to the in-context reversal curse using a controlled Father–Son relation probe in completion mode. Each example consists of a context sentence establishing a relationship, followed by a query prefix. In the Normal condition, the query is a copy-paste repetition of the relational statement up to the answer position (e.g., context: “The son of Michael is John.”; query: “The son of Michael is”). In the Reversed condition, the query inverts the relation direction (e.g., context: “The son of Michael is John.”; query: “The father of John is”). We evaluate the model distribution at the first answer position and report the negative log-likelihood (NLL; in nats) of two candidates: the target token (the correct answer) and the distractor token (the other name present in the prompt).
Figure 4 plots the resulting NLL for the target (green) and distractor (red) at the first predicted position, averaged over 1000 samples per condition. For readability, the y-axis is inverted so that higher (lower NLL) is better. Because lower NLL implies higher confidence, an error corresponds to the distractor having lower NLL than the target (red marker above green). In these figures, dashed lines are least-squares fits of the target NLL trend across the 12-50M token training runs (using the same checkpoints as in the data-scaling experiments). We also include pretrained GPT-2 and Mistral-7B as reference points to illustrate behavior at substantially larger pretraining scale.
@@
In the reversed condition, TG improves the probability of target token substantially faster than GPT-2 (target-NLL trend slope vs. ) and achieves lower target NLL in the reversed condition at 30M and 50M training dataset sizes. To summarize directional bias, we plot the reverse-direction margin where indicates a preference for the correct answer and indicates a distractor preference (a reversal error). In the reversed condition, improves monotonically for TG (approximately nats from 12M50M), indicating steadily reduced bias toward the distractor. In contrast, GPT-2’s margin becomes more negative through 30M and only slightly recovers at 50M, with a strong distractor preference persisting even for pretrained GPT-2 trained on web-scale data.This pattern suggests that for a standard token-stream Transformer, additional data may even amplify an order-bound shortcut (copying the wrong entity at a similar position) rather than capturing the underlying relationship. TG’s sentence-level contextualization reduces this shortcut pressure and moves the margin toward zero, with trends pointing toward a resolution of the in-context reversal curse (similar to that observed in Mistral) at a much lower pretraining scale.
Normal condition.
In the normal (copy-paste) completion, both models improve at nearly the same rate (slopes for GPT-2 vs. for TG). GPT-2 is more confident in absolute terms (lower target NLL), while TG shows a consistent offset that shrinks with additional training data. This is expected, as the normal condition represents a trivial in-context token sequence copying task; GPT-2 has direct access to the exact token sequence, while TG must infer the sequence from a compressed sentence state. Critically, TG matches GPT-2’s improvement trend in the normal condition while exhibiting a much higher improvement rate in the inverse condition. This supports our hypothesis that conditioning on fully contextualized latent states, rather than solely on position-tied token sequences, can mitigate the reliance on surface word order that causes the reversal curse.
@@
| In-context Sentence Memory | 30.2 | 19 | 85.6M |
@@ 4.4 Ablations
Table 1 reports ablations designed to isolate which TG components drive the gains observed in §4. All ablation models are trained on the same -token pretraining subset using the protocol in §4. For each run we report: (i) test perplexity (computed over lexical tokens only, matching §4); (ii) the number of non-embedding parameters ; and (iii) training throughput measured as sentence steps per second. Throughput is measured on a single NVIDIA A40 GPU with no model or data parallelism.
@@
The baseline TG stack alternates within-sentence self-attention and memory cross-attention layers. We evaluate higher-capacity variants in which each layer contains both mechanisms: a serial SelfCross ordering and a Parallel Self&Cross ordering. SelfCross improves test perplexity from , but increases model size from non-embedding parameters () and reduces throughput from sent./sec. The Parallel variant matches the same parameter and throughput cost but yields a smaller gain (29.7 PPL), suggesting that the serial ordering is a more effective way to scale TG under this design. Overall, these results show that allocating additional parameters to TG can yield further gains and that adding per-layer cross and self attention is an effective route to scale TG.
Gradient flow through memory is essential.
The most consequential ablation detaches sentence representations at memory write time, which substantially worsens test perplexity (); however, this increases the throughput () due to the reduced backpropagation depth. This sharp degradation indicates that TG’s gains do not come merely from introducing a recurrent state, but from training that state end-to-end: next-token losses on later sentences backpropagate through memory reads to optimize the parameters that produced earlier sentence gestalts.
External vs. in-context memory.
Placing sentence vectors in context as a prefix and relying on self-attention across the extended sequence (while preserving gradient flow through the memory vectors) largely retains TG’s modeling benefits (test PPL ), but is slower (throughput vs. sent./sec). This is consistent with the quadratic cost of self-attention: for sentence length and memory size , TG decouples processing into self-attention and cross-attention, while the in-context approach self-attends to the concatenated sequence in .
@@
The remaining ablations yield smaller but systematic degradations: extracting sentence representations from the last layer, removing context seeding, disabling EOS down-weighting, disabling the stream-length curriculum, and halving the maximum sentence length all increase test perplexity by – PPL (about –) relative to baseline. Among these, removing the stream curriculum produces the largest drop (PPL ), supporting the role of controlling backpropagation depth for stable optimization. Overall, among the components tested here, the dominant driver of performance is end-to-end training of sentence gestalts through differentiable memory, with the remaining components providing incremental improvements.

### v2

@@ 2.3 Training schedules
Sentence-stream curriculum
Retaining computation graphs through memory writes causes the effective backpropagation depth to grow with the number of sentences processed continuously (i.e., without resetting the memory between sentence streams), even though the number of sentence representations available in forward pass is capped at . Early in training, sentence representations are largely uninformative; long sentence streams therefore increase compute and optimization difficulty without providing useful long-range credit assignment. We address this with a curriculum over the maximum sentence-stream length used to slice training documents: we begin with shorter streams and periodically increase and re-chunk the training split (validation and test sets are never split into sentence streams). In our experiments with memory capacity , we start at sentences and increase by every 5 epochs. This expands the effective dependency span of optimization from 750 (average sentence length of ) to 2000 tokens over the course of training, while keeping early optimization focused on within-sentence token modeling and short-range memory use.
@@
In TG, end-of-sentence markers (<EOS>) are more frequent than lexical tokens, as they appear at the end of every sentence and are often comparatively easy to predict (strongly signaled by punctuation and syntax). To mitigate this frequency and easy-label imbalance, we apply a frequency-aware reweighting of the token-level loss that down-weights <EOS> targets after an initial warm-up epoch (weight 1.0 for epoch 1, then 0.05 thereafter) [33, 10]. This affects the training loss the model receives; separately, special-token targets (<EOS>/<EOD>) are excluded from reported perplexities and checkpoint selection, as described in §3.1.
@@ 3 Results
We evaluate the Thought Gestalt model through a series of experiments designed to measure data efficiency and representational robustness. First, we analyze TG’s scaling behavior by measuring how loss improves with training data and model size relative to a standard GPT-2 baseline, following the empirical scaling-law framework of Kaplan et al. [26]. Second, we benchmark TG’s data scaling against three architectural baselines that isolate key design concepts: (1) GPT-2 with Sentence Boundary Bias, which induces structural bias by retaining sentence start/end tokens (<BOS>, <EOS>) in the token stream; (2) Fixed Token-Span Recurrence, which replaces TG’s sentence-level processing with fixed-length token windows similar to prior recurrent work [7, 23]; and (3) GPT-2 + Gist Masking, an in-context gisting-style attention masking method [44], where tokens attend causally within a sentence and access preceding sentences only through compressed “gist” tokens that are computed in a single pass, without recurrence. Third, we assess relational direction generalization, which evaluates robustness to the in-context reversal curse given a controlled father–son probe [5, 34]. Finally, we conduct design ablations to identify how specific components, such as gradient flow through memory, drive performance gains.
3.1 Experimental Setup
@@ 3.2 Scaling Efficiency
We quantify the data and parameter efficiency of Thought Gestalt (TG) using the empirical scaling-law framework of Kaplan et al. [26]. Following this work, we measure the held-out test cross-entropy loss in nats/token and track model size as , the number of non-embedding parameters. When other factors are not the limiting bottleneck, Kaplan et al. observed approximate power-law relationships and , which appear as straight lines in log–log space. In this regime, architectural improvements typically manifest as a downward shift of the curve (lower loss at fixed resources). We first fit these laws to TG and a matched GPT-2 baseline trained under identical optimization and evaluation protocols described in §3.1 to compare their data and parameter efficiency.
@@
Figure 2(a) shows that TG achieves lower test perplexity at every dataset size. Fitting a power law to the test loss yields similar scaling exponents for the two models (TG: ; GPT-2: ), suggesting that over this range the gain is primarily due to an intercept shift. Across our tested range this corresponds to a – reduction in perplexity (e.g., at M tokens, TG reaches 23.2 test perplexity versus 24.0 for GPT-2; see Appendix E.1 for other perplexities). To express the improvement as data efficiency, we compute an effective data multiplier , where is TG’s achieved test loss at dataset size , and is the number of training tokens GPT-2 would need to reach loss according to the fitted GPT-2 scaling curve. We find –, meaning GPT-2 requires about 5–8% more training tokens to match TG’s loss; supporting the hypothesis that learning latent representations at more abstract levels than tokens can lead to more sample efficient learning.
@@ 3.3 Other Baselines
TG leverages sentence boundaries as a structural bias to process information in semantically coherent units. Here, we evaluate whether a standard transformer benefits simply from making this structure explicit in the token stream. To do this, we take the sentence-segmented output of the TG data pipeline (§2.1) and flatten it into a continuous sequence for a standard GPT-2 model. Crucially, we retain the <BOS> and <EOS> markers to explicitly delimit boundaries, while removing the padding tokens required for TG’s fixed-size sentence tensors. In this setup, <BOS> predicts the sentence’s first lexical token, the last lexical token predicts <EOS>, and <EOS> predicts the <BOS> of the subsequent sentence. The model architecture and attention mechanism remain identical to the GPT-2 baseline. To avoid making the comparison hinge on predicting frequent boundary events, we compute validation loss for checkpoint selection and report perplexity over lexical tokens only, excluding special-token label positions, matching the TG evaluation protocol (§3.1). Figure 4(left) shows that inducing sentence boundary bias improves GPT-2 performance across all dataset sizes, with the most significant gains in the low-data regime (test PPL decreases at 20M; see Table 3). However, the improvement shrinks with scale (e.g., at 50M), and TG surpasses this baseline once more data is available (TG: vs. at 50M). This indicates that boundary tokens alone provide a helpful structural inductive bias for token LMs, but this is not sufficient to replicate the efficiency gains derived from TG’s learning and reuse of contextualized sentence-level latent states.
@@
TG’s attention pattern—full causal token interaction within a sentence, and access to earlier content only through compact sentence vectors—is similar to gisting: tokens attend to their current segment and can only access prior segments through a small set of “gist” representations learned via attention masking in a standard transformer [44]. To test whether this attention distribution mechanism can reproduce TG’s gains without recurrence or an external memory, we implement a GPT-2 baseline that (i) inserts <BOS>/<EOS> boundaries in the token stream and (ii) applies an additive attention-bias mask that restricts each token to attend causally within its own sentence, while accessing previous sentences only via each sentence’s last token, <EOS>, acting as a “gist” token. This matches TG’s attention connectivity, but removes recurrence, which has a key stabilizing effect: TG exposes fully contextualized sentence vectors to every layer via cross-attention, whereas GPT-2 + gist masking forces layers to read from in-context compression tokens whose representations are formed in parallel with all other tokens. In the original work, a prompt prefix is compressed into learned gist tokens which are computed in a separate pass for caching (a one iteration recurrence); here we compute all sentence gists in a single pass, without caching, to ablate recurrence. Empirically, the gist-masking baseline performs worse across all data scales (test PPL 31.9 at 50M; Table 5), far behind both GPT-2 (24.0 at 50M) and TG (23.2 at 50M), as seen in Figure 4 (right). These results show that in-context compression, restricting direct access to prior token, can degrade pretraining when compressed representations are formed concurrently with the later tokens that depend on them, without recurrence.
@@
Prior work on the reversal curse characterizes a training-time directional asymmetry: models trained (or fine-tuned) on relational statements in one direction (“A is B”) often fail to generalize to the inverse (“B is A”) [5, 34]. Here we evaluate a distinct, inference-time phenomenon: an in-context reversal curse, where the relation is stated in the prompt, and the completion query asks for either the same direction or its inverse. Berglund et al. [5] note that reversal curse is not observed when relations are expressed in-context for the large models they evaluate (including GPT-3.5/4) [46]; we find that a substantial in-context directional bias persists at smaller scales, including the pretrained Hugging Face GPT-2 XL checkpoint (1.5B parameters; WebText pretraining on 40GB of text, roughly 9 billion tokens) [45]. Furthermore, we show that this in-context reversal error is resolved only at the scale of Mistral 7B (7 billion parameters, trained on trillions of tokens) [25]. To our knowledge, prior work has not documented the in-context directional asymmetry of smaller scale models.
Our in-context reversal curse experiments use Father–Son relation probes in completion mode. Each example consists of a sentence establishing a relationship, followed by a query prefix. In the Normal condition, the query repeats the relational statement up to the answer position (e.g., context: “The son of Michael is John.”; query: “The son of Michael is”). In the Reversed condition, the query inverts the relation direction (e.g., context: “The son of Michael is John.”; query: “The father of John is”). We evaluate the model token-probability distribution at the first answer position and report the negative log-likelihood (NLL; in nats) of two candidates: the target token (correct answer) and the distractor token (other name present in the prompt). Figure 5 visualizes the results. In the scatter plots (Fig. 5, top-left and bottom row), the y-axis is inverted so that higher marker placement (lower NLL) indicates higher token probability (higher model confidence). An in-context reversal error corresponds to the distractor having lower NLL than the target (a red marker above the corresponding green marker), equivalently a negative reverse-direction margin (Fig. 5, top-right). Dashed lines are least-squares fits of the target-NLL trend across the dataset-scaling checkpoints (the same runs as in Fig. 2a).
@@
In the reversed condition, TG improves the probability of target token substantially faster than GPT-2 (Fig. 5, top-left; target-NLL trend slope vs. ) and achieves lower target NLL in the reversed condition at 30M and 50M training dataset sizes. To summarize directional bias, we plot the reverse-direction margin where indicates a preference for the correct answer and indicates a distractor preference (a reversal error; Fig. 5, top-right). In the reversed condition, improves monotonically for TG (approximately nats from 12M50M), indicating steadily reduced bias toward the distractor. In contrast, GPT-2’s margin becomes more negative through 30M and only slightly recovers at 50M, with a strong distractor preference persisting even for the pretrained GPT-2 XL checkpoint (Fig. 5, bottom-left). This pattern suggests that for a standard token-stream transformer, additional data may amplify an order-bound shortcut (copying the wrong entity at a similar position) rather than capturing the underlying relationship. TG’s sentence-level contextualization reduces this shortcut pressure and moves the margin toward zero, with trends pointing toward a resolution of the in-context reversal curse at a much lower pretraining scale than GPT-2.
Figure 5 (bottom-left) shows that the pretrained GPT-2 XL checkpoint succeeds in the Normal (copy-paste) query (target preferred over distractor) but exhibits a large in-context reversal error in the Reversed query (distractor preferred over target). In contrast, Mistral 7B (a 7 billion parameter foundation model likely pretrained on order of trillions of tokens [12]) prefers the target in both directions, indicating that this in-context directional bias is largely absent at 7B scale [25].
TG vs. GPT-2: Normal condition.
In the normal (copy-paste) completion, both models improve at nearly the same rate (Fig. 5, bottom panels; slopes for GPT-2 vs. for TG). GPT-2 is more confident in absolute terms (lower target NLL), while TG shows a consistent offset that shrinks with additional training data. This is expected, as the normal condition represents a trivial in-context token-sequence copying task; GPT-2 has direct access to the exact token sequence, while TG must infer the sequence from a compressed sentence state. Critically, TG matches GPT-2’s improvement trend in the normal condition while exhibiting a much higher improvement rate in the inverse condition (Fig. 5, top-left and top-right). This supports our hypothesis that conditioning on fully contextualized latent states, rather than solely on position-tied token sequences, can mitigate the reliance on surface word order that causes the reversal curse.
@@
| In-context working memory | 30.2 | 19 | 85.6M |
@@ 3.5 Ablations
Table 1 reports ablations designed to isolate which TG components drive the gains observed in §3. All ablation models are trained on the same -token pretraining subset using the protocol in §3.1. For each run we report: (i) test perplexity (computed over lexical tokens only, matching §3.1); (ii) the number of non-embedding parameters ; and (iii) training throughput measured as sentence steps per second. Throughput is measured on a single NVIDIA A40 GPU with no parallelization.
@@
The baseline TG stack alternates within-sentence self-attention and memory cross-attention layers. We evaluate higher-capacity variants in which each layer contains both attention mechanisms before the feedforward network: a serial SelfCross attention ordering and a parallel Self&Cross ordering. SelfCross improves test perplexity from , but increases model size from non-embedding parameters () and reduces throughput from sent./sec. The Parallel variant matches the same parameter and throughput cost but yields a smaller gain (29.7 PPL), suggesting that the serial ordering is a more effective way to scale TG under this design. Overall, these results show that allocating additional parameters to TG can yield further gains and that adding per-layer cross and self attention is an effective route to scale TG.
Working memory and gradient flow through memory are essential.
The largest performance drop comes from removing the working memory entirely by replacing every cross-attention block with within-sentence causal self-attention (Table 1). This ablation effectively reduces TG into a sentence-local decoder whose token context resets at every sentence boundary (effectively a GPT-2 with a one-sentence context window), and test perplexity collapses (). This sharp degradation indicates that TG’s performance fundamentally relies on conditioning token generation on earlier context via sentence gestalts, consistent with the broader evidence that TG increasingly up-weights and routes learning signal through the working memory pathway as training progresses (Appendices C and D). The next most consequential ablation detaches the gradients of sentence representations when updating the working memory, which considerably worsens test perplexity (); despite increasing the throughput modestly () due to the reduced backpropagation depth. This degradation indicates that TG’s gains do not come merely from introducing a recurrent state, but from training that latent state end-to-end: next-token losses on later sentences backpropagate through memory reads to optimize the parameters that produced earlier sentence gestalts.
External vs. in-context working memory.
Placing sentence vectors in context as a prefix and relying on self-attention across the extended sequence (while preserving gradient flow through the memory vectors) largely retains TG’s modeling benefits (test PPL ), but is slower (throughput vs. sent./sec). This is consistent with the quadratic cost of self-attention: for token sentence length and memory size , TG decouples processing into self-attention and cross-attention, while the in-context approach self-attends to the concatenated sequence in .
@@
The remaining ablations, such as the in-context memory ablation, yield modest but interpretable degradations: extracting sentence representations from the last layer, removing context seeding, disabling EOS down-weighting, disabling the stream-length curriculum, and halving the maximum sentence length all increase test perplexity by – PPL (about –) relative to baseline. Among these, removing the stream curriculum produces the largest drop (PPL ), supporting the role of controlling backpropagation depth for stable optimization. Overall, among the components tested here, the dominant driver of performance is end-to-end training of sentence gestalts through differentiable memory, with the remaining components providing incremental improvements.
4 Related Works
Learning sentence representations.
Large Concept Model (LCM) [56] similarly operates at the sentence level, but trains an autoregressive model to predict the next sentence embedding in a frozen sentence representation space (SONAR), which is then decoded into tokens, rather than directly predicting a token sequence. In contrast, the Thought Gestalt (TG) model does not use a frozen encoder or any separate sentence-level loss: token and sentence spaces are learned jointly using the same next-token prediction objective. This avoids brittle auxiliary losses while still yielding contextualized sentence embeddings that can be reused as high-level states representing prior content.
Sequence compression and gisting.
In TG, sentence representations play the role of compression tokens: the model extracts a single gist vector per sentence from a mid–upper layer after full contextualization with within-sentence tokens and the sentence memory, and writes it into the working memory without detaching its computation graph. The only supervision for these sentence compression tokens comes from the same next-token losses on later sentences, backpropagated through the memory. Unlike Compressive transformers, TG does not introduce a separate compression loss; and unlike standard gisting, the same fully contextualized sentence representation is reused in the cross-attention path at every layer (each transformer layer applies its own learned projections to these vectors). Thus, even early layers can query a stable, contextualized summary of prior content rather than raw, noisy activations, helping alleviate long-range contextualization errors [32].
Recurrence and memory in transformers.
Another design axis concerns extending transformers to reuse past representations to enable accessing content beyond a fixed context window. transformer-XL caches hidden states from previous segments and lets the current segment attend to them, providing segment-level recurrence but truncating gradients at the cached states [11]. Block-Recurrent transformers apply a transformer layer recurrently over blocks of tokens, with a recurrent state that each layer attends to, combining local self-attention within a block with an RNN-like state that carries information across blocks [23]. Recurrent Memory transformer adds dedicated memory tokens that are passed between segments and updated by self-attention, providing a differentiable memory channel across long sequences [7]. In a complementary direction, Latent Thought Models (LTMs) introduce a memory of sequence-specific latent “thought” vectors inferred via iterative variational Bayes updates (inference-time computation) to guide a transformer decoder [29]. Memorizing transformers, on the other hand, add an external key–value store of past activations that can be queried with kNN retrieval at inference time, extending the effective context while keeping the memory non-differentiable [58].
TG shares with these models the design principles of reusing past computation, but its recurrent state resides in an external differentiable memory of semantically coherent sentence gestalts rather than token-level caches or memory tokens encoding arbitrary token sequences, and these gestalts are trained only via a robust language modeling objective: the next-token prediction loss. This design ties recurrence, compression, and sentence structure together: long-range information is carried forward in compact sentence representations that are optimized end-to-end for generative performance. These semantically grounded units are natural candidates for storage and retrieval in a long-term memory system without the need to increase model parameters, in the spirit of the Memorizing transformer but using sentence- or thought-aligned chunks that are more effective than arbitrary token blocks [30, 6].